OpenStudy (anonymous):

Solve Sqrt x+14-12=x.

OpenStudy (help!!!!):

..

OpenStudy (anonymous):

\[\sqrt{x+14}-12=x\] add 12 get \[\sqrt{x+14}=x+12\] then square both sides \[x+14=(x+12)^2=x^2+24x+144\] and now you have a quadratic equation to solve

OpenStudy (anonymous):

\[x^2+23x+134=0\] and so on

OpenStudy (anonymous):

don't forget to check both answers as one may not work

OpenStudy (anonymous):

I don't get it -.-

OpenStudy (anonymous):

all steps are written out, except for solving the quadratic equation first isolate the radical on one side of the equation then square both sides to remove the radical finally solve for \(x\)

OpenStudy (hba):

@bunny89 Is the Sqrt on the whole thing ?